# Node-like as a Whole: Structure-aware Searching and Coarsening for Graph Classification

### v2

@@ Abstract.
Graph Transformers (GTs) have made remarkable achievements in graph-level tasks. However, most existing works regard graph structures as a form of guidance or bias for enhancing node representations, which focuses on node-central perspectives and lacks explicit representations of edges and structures. One natural question is, can we treat graph structures node-like as a whole to learn high-level features? Through experimental analysis, we explore the feasibility of this assumption. Based on our findings, we propose a novel multi-view graph representation learning model via structure-aware searching and coarsening (GRLsc) on GT architecture for graph classification. Specifically, we build three unique views, original, coarsening, and conversion, to learn a thorough structural representation. We compress loops and cliques via hierarchical heuristic graph coarsening and restrict them with well-designed constraints, which builds the coarsening view to learn high-level interactions between structures. We also introduce line graphs for edge embeddings and switch to edge-central perspective to construct the conversion view. Experiments on eight real-world datasets demonstrate the improvements of GRLsc over 28 baselines from various architectures.
Graph Classification, Graph Coarsening, Graph Transformer
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data miningccs: Theory of computation Graph algorithms analysisccs: Computing methodologies Supervised learning
1. Introduction
@@
Graph Neural Networks (GNNs) have become a significant approach to Graph Representation Learning (GRL) recently, achieving remarkable performance on various node-level (such as node classification (Chen et al., 2023a) and link prediction (Chamberlain et al., 2023)) and graph-level (such as graph classification (Niepert et al., 2016; Xinyi and Chen, 2019; Yao et al., 2023)) tasks. GNNs model the graph structure implicitly, leveraging graph topology information to aggregate neighborhood node features. However, Xu et al. (2019) have proved that the aggregation mechanism of GNNs has local limitations, which has encouraged more researchers to move from GNNs to GTs (Liu et al., 2023a; Chen et al., 2023b; Wu et al., 2023). GTs utilize position encoding to represent the graph topology globally, jumping out of the neighborhood restriction and interacting between distant node pairs. Although GTs can learn long-distance structural information, most GNNs and GTs use graph structures as guidance (Huo et al., 2023) or bias (Ying et al., 2021) to obtain better node representations rather than directly representing them.
Structural information is crucial for graph-level tasks, especially Graph Structural Learning (GSL) (Zhang et al., 2023c; Zou et al., 2023; Zhang et al., 2023a). GSL aims to learn the optimized graph structure and representation from the original graph, further improving the model performance on downstream tasks (e.g. graph classification). Kernel-based methods (Shervashidze et al., 2009, 2011) lead the way of GSL in the early stage, which selects representative graph structures carefully to test the isomorphism and update weights in classifiers. However, this category of methods suffers due to computational bottlenecks of the pair-wise similarity (Wu et al., 2021; Yang et al., 2023). With the development of deep learning, researchers have proposed lots of new graph learning paradigms such as Graph Contrastive Learning (GCL) (Sun et al., 2020; Ma et al., 2023; Liu et al., 2023b). Using various data augmentation strategies like node dropping (Ma et al., 2023), edge perturbation (Zhang et al., 2023b), attribute masking (Yuan et al., 2023), and subgraph sampling (Liu et al., 2023b), GCL ensures the semantic matching between views and enhances the robustness of the model against multifaceted noises. Though GCL has made considerable achievements, views obtained through the strategies above keep the original structures, which restricts model capabilities since limited usage of high-level structural information.
To fully utilize topological information and capture high-level structures, one natural question arises as to whether we can treat graph structures node-like as a whole for graph-level tasks. Take the molecule graph as an example, specific structures such as molecular fragments and functional groups contain rich semantics. Random perturbation of these structures will produce additional structural information, which has been proved to be invalid nevertheless (Zhu et al., 2023). Some researchers have focused on this problem and defined the overall treatment idea as Graph Compression (Liu et al., 2022), Pooling (Wei et al., 2021; Baek et al., 2021; Lv et al., 2023), and Coarsening (Zhang et al., 2023b). Inspired by Convolutional Neural Networks (CNNs) to compress the set of vectors into a compact representation (Wu et al., 2022a), this series of works aims to address the flat architecture of GNNs, emphasizing the hierarchical structures of graphs (Ying et al., 2018).
@@
Figure 1. Pre-experiment on a toy molecule graph set, which shows four graphs with different features, all containing a benzene ring. (B) illustrates the pairwise comparison between four graphs. (C) presents the comparison between the graph and its coarsening view. We leverage two measurements of Euclidean Distance (bar plots, lower for closer) and Cosine Similarity (line plots, higher for closer) to reveal the latent relations between graphs.
@@
We further explore the feasibility of treating graph structures node-like as a whole through a set of pre-experiments. Figure 1 shows the results. The pre-experiments follow the plain message-passing mechanism and summation readout function. We train the pre-experimental model on four toy graphs and obtain representations of the whole graph, benzene ring structure, and main distinguishing atom sets. We measure the spatial distance and angular separation between representations by Euclidean Distance and Cosine Similarity, respectively. Figure 1(B) shows the comparison between different graphs. We divide the graph pairs into three groups according to different atom positions and different atom numbers. They are: (1) same position, different numbers: AC and BD; (2) different positions, same number: AB and CD; (3) different positions, different numbers: AD and BC. The experiments present instructive results, where the gap between loop is less, w.r.t. graph and diff. It highlights that the prime factor causing the differences lies behind the main distinguish atom sets rather than the benzene ring. Thus, we compress the benzene ring into one node with the same settings and obtain the coarsening view of each graph. Figure 1(C) shows the distance between the original graph and its coarsening view. Pairs such as A and C, whose main distinguish atom sets are arranged adjacent, leave a relatively closer cosine similarity, while B and D show a significant difference. It indicates that simple coarsening will lose part of structural information. In summary, pre-experiments tell us two enlightenments: (1) some structures contribute relatively less to distinguish graphs, which can further turn into a coarsening view to magnify the high-level structural information; (2) after treating structures node-like as a whole, some of the structural information may be traceless, calling for additional consideration for the relative position of neighborhoods.
Next, we reinforce the conclusions drawn from the pre-experiment with other real-world scenarios. Many systems, like recommendation (Yan et al., 2024), education (Rehm et al., 2024), and e-commerce (Fan et al., 2023), reveal imbalanced distributions of nodes, edges, and attributes, forming specific structures noteworthy. For example, cliques are groups of nodes with a high internal closeness and low external (Tsitsulin et al., 2023). From the node-central view, it is necessary to analyze the internal connections. Rich connections provide sufficient information for node-level analysis, such as Meanwhile, standing on the system side, the sparse edges outside cliques may generate high-level structural information, and many works (Xu et al., 2019; Ying et al., 2018) have proved that models can benefit from them. Low-density edges outside cliques represent how two cliques connect. It can answer some interesting questions, such as: what other sports will the group interested in football be interested in, which goods will the group who buy more electronic products buy, and so on.
Therefore, based on the above discovery, we focus on loops and cliques and introduce two views: a coarsening view and a conversion view to emphasize individual components of the graph. The former coarsens graph structures into one node based on clustering to learn high-level structural information. The latter transforms nodes and edges from the node-central perspective to edge-central, which highlights relative position and is complementary to the loss of coarsening. Specifically, we construct the graph coarsening view through a heuristic algorithm restricted by well-designed constraints and build the line graph conversion view to augment relative position information. Together with the original graph view, we train separate GT for each view for graph encoding. Finally, we concatenate multi-view representation vectors as the final embedding of an entire graph.
@@
  • We design a novel multi-view graph representation learning model via structure-aware searching and coarsening (GRLsc), which leverages three views to learn more comprehensive structural information for downstream graph classification tasks.
@@
  • We propose a hierarchical heuristic algorithm to compress the loops and cliques and construct the graph coarsening view, which captures high-level topological structures.
  • We introduce the line graph conversion view, which retains relative position information between neighbor nodes.
@@
  • We verify the performance of GRLsc on 8 datasets from 3 categories, compared with 28 baselines from 6 categories. GRLsc achieves better results than SOTAs.
2. Related Work
2.1. Graph Pooling.
@@ 2. Related Work
Graph Pooling (Wei et al., 2021; Baek et al., 2021; Lv et al., 2023) is one of the pivotal techniques for graph representation learning, aiming to compress the input graph into a smaller one. Due to the hierarchy characteristics of graphs, hierarchical pooling methods have become mainstream nowadays. Methods fall into two categories depending on whether breaking nodes. The first category conducts new nodes via graph convolution. The pioneer in hierarchical pooling, DIFFPOOL (Ying et al., 2018), aggregates nodes layer by layer to learn high-level representations. Inspired by the Computer Vision field, CGKS (Zhang et al., 2023b) constructs the pooling pyramid under the contrastive learning framework, extending the Laplacian Eignmaps with negative sampling. MPool (Islam et al., 2023) takes advantage of motifs to capture high-level structures based on motif adjacency. The second category condenses graphs by sampling from old nodes. SAGPool (Lee et al., 2019) masks nodes based on top-rank selection to obtain the attention mask subgraphs. CoGSL (Liu et al., 2022) extracts views by perturbation, reducing the mutual information for compression. MVPool (Zhang et al., 2023a) reranks nodes across different views with the collaboration of attention mechanism, then selects particular node sets to preserve underlying topological.
2.2. Graph Coarsening and Condensation.
@@
Similar to Graph Pooling, both Graph Coarsening (Pang et al., 2021; Chen et al., 2023c; Zhang and Zhao, 2024) and Graph Condensation (Jin et al., 2022b, a) are graph reduction methods that simplify graphs while preserving essential characteristics (Hashemi et al., 2024). Generally speaking, Graph Coarsening groups and clusters nodes into super node sets using specified aggregation algorithms. It preserves specific graph properties considered high-level structural information, revealing hierarchical views of the original graph. L19 (Loukas, 2019) proposes a restricted spectral approximation with relative eigenvalue error. SS (Zhang and Zhao, 2024) conducts both intra- and inter-cluster features to strengthen the fractal structures discrimination power of models. In addition to the above spectral methods, researchers try to find other ways to make measurements. KGC (Chen et al., 2023c) calculates graphs equipped with Gromov-Wasserstein distance.
@@
Graph Condensation, first introduced in (Jin et al., 2022b), leverages learnable downstream tasks’ information to minimize the loss between the original graph and some synthetic graphs. They propose a method based on gradient matching techniques, called GCond (Jin et al., 2022b), to condense the structures via MLP. However, GCond involves a nested loop in steps. Additional requirements for scalability led to the creation of DosCond (Jin et al., 2022a), EXGC (Fang et al., 2024), and many other works (Li et al., 2023; Zhang et al., 2024). DosCond (Jin et al., 2022a) is the first work focusing on graph classification tasks via graph condensation. EXGC (Fang et al., 2024), on the other hand, heuristically identifies two problems affecting graph condensation based on gradient matching and proposed solutions. Both types of work hope to achieve the same downstream task effect as the original graph while reducing the graph scale. For a further detailed understanding of those techniques, we recommend reading the survey (Hashemi et al., 2024).
2.3. Graph Transformers.
@@
Graph Transformers (GTs) (Chen et al., 2023b; Wu et al., 2023) alleviate the problems of over-smoothing and local limitations of GNNs, which has attracted the attention of researchers. The self-attention mechanism learns long-distance interactions between every node pair and shows tremendous potential for various scenarios. Graphormer (Ying et al., 2021) integrates edges and structures into Transformers as biases, outperforming in graph-level prediction tasks. U2GNN (Nguyen et al., 2022) alternates aggregation function leveraging Transformer architecture. Exphormer (Shirzad et al., 2023) builds an expansive GT leveraging a sparse attention mechanism with virtual node generation and graph expansion. In addition, due to the quadratic computational constraint $O(N^{2})$ of self-attention, several works (Wu et al., 2022b; Liu et al., 2023a) have been proposed to focus on a scalable and efficient Transformer. The former puts forward a new propagation strategy adapting arbitrary nodes in scale. The latter combines pooling blocks before multi-head attention to shrink the size of fully connected layers. To summarize, we design our model based on the capabilities of GTs, which can effectively aggregate features of distant nodes using attention mechanisms and overcome the limitations of local neighborhoods.
Figure 2. Framework of our proposed model GRLsc (center). We show an example input graph $G_{i}$ and the corresponding intermediate output coarsening graph (bottom right). We will explain the coarsening and conversion details later.
3. Preliminaries
3.1. Notations
3.1.1. Graphs.
@@ 3.1. Notations
3.2. Universal Graph Transformer (U2GNN)
U2GNN (Nguyen et al., 2022) is a GNN-based model that follows the essential aggregation and readout pooling functions. Xu et al. (2019) claim that a well-designed aggregation function can further improve the performance. Thus, U2GNN replaces the plain strategy which GNN uses with the following aggregation function:
(1) $h_{t,v}^{(k)}=\mathbf{AGG}_{\mathsf{att}}(h_{t-1,v}^{(k)}),$
@@ 3.2. Universal Graph Transformer (U2GNN)
4. Methodology
4.1. Overview
Figure 2 shows the framework diagram of our proposed model GRLsc, which mainly contains three parts: multi-view construction of graphs, graph encoders, and classifier (omitted in the figure). Given an input graph $G_{i}$, GRLsc constructs three views step by step. Keep the original graph as $G_{o}$, and coarsen $G_{i}$ via the Loop and Clique Coarsening (LCC) Block to build view $G_{lcc}$. After that, we design the Line Graph Conversion (LGC) Block to transfer $G_{lcc}$ to view $G_{lgc}$. Subsequently, GRLsc trains a GT for each view separately to obtain graph-level representation, which is directly concatenated and fed into the downstream classifier.
This chapter is structured as follows: Section 4.2 and Section 4.3 describe LCC Block and LGC Block, respectively. Section 4.4 explains the rest of the details of GRLsc.
4.2. Loop and Clique Coarsening
4.2.1. Algorithm Description
Algorithm 1 Loop and Clique Coarsening (LCC) Block

### v5

# An Efficient Loop and Clique Coarsening Algorithm for Graph Classification

@@ Abstract
Graph Transformers (GTs) have made remarkable achievements in graph-level tasks. However, most existing works regard graph structures as a form of guidance or bias for enhancing node representations, which focuses on node-central perspectives and lacks explicit representations of edges and structures. One natural question arises as to whether we can leverage a hypernode to represent some structures. Through experimental analysis, we explore the feasibility of this assumption. Based on our findings, we propose an efficient Loop and Clique Coarsening algorithm with linear complexity for Graph Classification (LCC4GC) on GT architecture. Specifically, we build three unique views, original, coarsening, and conversion, to learn a thorough structural representation. We compress loops and cliques via hierarchical heuristic graph coarsening and restrict them with well-designed constraints, which builds the coarsening view to learn high-level interactions between structures. We also introduce line graphs for edge embeddings and switch to edge-central perspective to alleviate the impact of coarsening reduction. Experiments on eight real-world datasets demonstrate the improvements of LCC4GC over 31 baselines from various architectures. Our code is available: https://github.com/NickSkyyy/LCC-for-GC.
keywords:
Graph Representation Learning , Graph Coarsening , Graph Classification
journal: Pattern Recognition
[…]
[l1]organization=College of Computer Science, Nankai University, country=China
1 Introduction
@@
Graph Transformers (GTs) have succeeded in graph classification tasks [1, 2, 3]. GTs utilize position encoding to represent the graph topology globally, jumping out of the neighborhood restriction and interacting between distant node pairs. Although GTs can learn long-distance structural information, most GNNs and GTs use graph structures as guidance [4] or bias [5] to obtain better node representations rather than directly representing them.
Structural information is crucial for graph-level tasks, urging Graph Structural Learning (GSL) [6, 7, 8]. GSL aims to learn the optimized graph structure and representation from the original graph, further improving the model performance on downstream tasks (e.g. graph classification). To fully utilize topological information, one natural question arises as to whether we can leverage a hypernode to represent some structures. Take the molecule graph as an example, specific structures such as molecular fragments and functional groups contain rich semantics. When we coarsen these structures into hypernodes, the new graph contains the interaction information with molecule fragments and functional groups as basic units, called high-level structure. Research [9] has proved that models can benefit from them. Meanwhile, random perturbation of these structures will produce additional structural information, which has been proved to be invalid nevertheless [10].
@@
Figure 1: Pre-experiment on a toy molecule graph set, which shows four graphs with different features, all containing a benzene ring. (B) illustrates the pairwise comparison between four graphs. (C) presents the comparison between the graph and its coarsening view. We leverage two measurements of Euclidean Distance (bar plots, lower for closer) and Cosine Similarity (line plots, higher for closer) to reveal the latent relations between graphs.
@@
We further explore the feasibility of representing structures via hypernodes through a set of pre-experiments. Figure 1 shows the results. We train a plain GCN and use summation as the readout function, obtaining three representations: the benzene ring (loop), main distinguishing atom sets (diff), and whole graph (graph). We measure the spatial distance and angular separation between representations by Euclidean Distance and Cosine Similarity, respectively. Figure 1(B) shows the comparison between different graphs. The experiments present instructive results, where the gap between loop is less, w.r.t. graph and diff. It highlights that the prime factor causing the differences lies behind the main distinguish atom sets rather than the benzene ring. Thus, we coarsen the benzene ring into one node with the same settings and obtain the coarsening view of each graph. Figure 1(C) shows the distance between the original graph and its coarsening view. Pairs such as A and C, whose main distinguish atom sets are arranged adjacent, leave a relatively closer cosine similarity, while B and D show a significant difference. It indicates that simple coarsening will lose part of structural information [11]. In summary, pre-experiments tell us two enlightenments: (1) some structures contribute relatively less to distinguish graphs, which can further turn into a coarsening view to magnify the high-level structural information; (2) after coarsening these structures into hypernodes, some of the structural information may be traceless, calling for additional consideration for the relative position of neighborhoods.
Some researchers have focused on this problem and defined the overall treatment idea as Graph Compression [12], Pooling [13, 14, 15], and Coarsening [16, 17]. This series of works aims to compress the set of nodes into a compact representation [18] or emphasize the hierarchical structures of graphs [9]. Despite all achievements, there are still many challenges. (1) Unsuitability coarsening. Some works aim to achieve better results on different tasks, such as graph reconstruction [19] and node classification [20], where the coarsened perspective is not necessarily helpful for graph classification. (2) Poor efficiency. Identifying the whole set of structures (e.g., motifs [21]) has an unbearable computational overhead, especially for those condensing methods [20, 22]. (3) Distribution shift. Previous studies have shown that the model will exhibit poor generalization when there is a difference between the training and testing set distribution [23]. A deeper coarsening layer reduces abundant graph elements, lacking an effective alignment of the original graph, which has poor interpretability [11]. It further hinders the credibility of the coarsening graph view.
Therefore, we propose an efficient coarsening algorithm based on the above findings and challenges. Our model relies on the GT framework and aims to achieve better performance on graph classification tasks (challenge 1). To address the issue posed by challenge 2, we focus on coarsening two widespread classes of structures, loops, and cliques as a trade-off and implement a linear algorithm with well-designed constraints. Finally, we solve challenge 3 by introducing line graph conversion to construct a multi-view learning framework. The multi-view fusion can preserve the original graph distribution for learning. On the other hand, a shallow coarsening layer reveals clearer correspondences and provides more reliable interpretation guarantees.
@@
  • 1.
@@
    We propose an efficient Loop and Clique Coarsening algorithm with linear complexity for Graph Classification (LCC4GC), which captures high-level topological structures.
  • 2.
    We introduce line graph conversion to conduct multi-view graph representation learning on the GT framework, which learns more comprehensive and reliable structural information.
  • 3.
@@
    We verify the performance of LCC4GC on graph classification tasks, compared with 28 baselines from six categories under eight datasets from three categories. LCC4GC achieves better results in most cases.
2 Related Work
2.1 Graph Pooling
@@
Graph Pooling [13, 14, 15] is one of the pivotal techniques for graph representation learning, aiming to compress the input graph into a smaller one. Due to the hierarchy characteristics of graphs, hierarchical pooling methods have become mainstream nowadays. Methods fall into two categories depending on whether breaking nodes. The first category conducts new nodes via graph convolution. The pioneer in hierarchical pooling, DIFFPOOL [9], aggregates nodes layer by layer to learn high-level representations. Inspired by the Computer Vision field, CGKS [16] constructs the pooling pyramid under the contrastive learning framework, extending the Laplacian Eignmaps with negative sampling. MPool [24] takes advantage of motifs to capture high-level structures based on motif adjacency. The second category condenses graphs by sampling from old nodes. SAGPool [25] masks nodes based on top-rank selection to obtain the attention mask subgraphs. CoGSL [12] extracts views by perturbation, reducing the mutual information for compression. MVPool [26] reranks nodes across different views with the collaboration of attention mechanism, then selects particular node sets to preserve underlying topological.
2.2 Graph Coarsening and Condensation
@@
Similar to Graph Pooling, both Graph Coarsening [27, 28, 29] and Graph Condensation [20, 22] are graph reduction methods that simplify graphs while preserving essential characteristics [11]. Generally speaking, Graph Coarsening groups and clusters nodes into super node sets using specified aggregation algorithms. It preserves specific graph properties considered high-level structural information, revealing hierarchical views of the original graph. L19 [30] proposes a restricted spectral approximation with relative eigenvalue error. SS [29] conducts both intra- and inter-cluster features to strengthen the fractal structures discrimination power of models. In addition to the above spectral methods, researchers try to find other ways to make measurements. KGC [28] calculates graphs equipped with Gromov-Wasserstein distance.
@@
Graph Condensation, first introduced in [20], leverages learnable downstream tasks’ information to minimize the loss between the original graph and some synthetic graphs. They propose a method based on gradient matching techniques, called GCond [20], to condense the structures via MLP. However, GCond involves a nested loop in steps. Additional requirements for scalability led to the creation of DosCond [22], EXGC [31], and many other works [32, 33]. DosCond [22] is the first work focusing on graph classification tasks via graph condensation. EXGC [31], on the other hand, heuristically identifies two problems affecting graph condensation based on gradient matching and proposed solutions. Both types of work hope to achieve the same downstream task effect as the original graph while reducing the graph scale. For a further detailed understanding of those techniques, we recommend reading the survey [11].
2.3 Graph Transformers
@@
Graph Transformers (GTs) [34, 35, 36] alleviate the problems of over-smoothing and local limitations of GNNs, which has attracted the attention of researchers. The self-attention mechanism learns long-distance interactions between every node pair and shows tremendous potential for various scenarios. Graphormer [5] integrates edges and structures into Transformers as biases, outperforming in graph-level prediction tasks. U2GNN [37] alternates aggregation function leveraging Transformer architecture. Exphormer [38] builds an expansive GT leveraging a sparse attention mechanism with virtual node generation and graph expansion. In addition, due to the quadratic computational constraint $O(N^{2})$ of self-attention, several works [39, 40] have been proposed to focus on a scalable and efficient Transformer. The former puts forward a new propagation strategy adapting arbitrary nodes in scale. The latter combines pooling blocks before multi-head attention to shrink the size of fully connected layers. To summarize, we design our model based on the capabilities of GTs, which can effectively aggregate features of distant nodes using attention mechanisms and overcome the limitations of local neighborhoods.
3 Preliminaries
Graphs
@@ 3 Preliminaries
Universal Graph Transformer (U2GNN)
U2GNN [37] is one of the SOTAs in graph classification tasks, a GNN-based model that follows the essential aggregation and readout pooling functions. Xu et al. [41] claim that a well-designed aggregation function can further improve the performance. Thus, U2GNN replaces the plain strategy which GNN uses with the following aggregation function:
$h_{t,v}^{(k)}=\mathbf{AGG}_{\mathsf{att}}(h_{t-1,v}^{(k)}),$ (1)
@@
4 Methodology
4.1 Overview
Figure 2: Overview of our work. (A) The framework of LCC4GC. LCC contains two procedures: (B) clique coarsening and (C) loop coarsening. We design two constraints, hierarchy depth, and loop length, to control the graph coarsening process. (D) An example of how LCC and LGC work.
Figure 2(A) shows the framework diagram of our proposed model LCC4GC, which mainly contains three parts: multi-view construction of graphs, graph encoders, and classifier (omitted in the figure). Given an input graph $G_{i}$, LCC4GC constructs three views step by step. Keep the original graph as $G_{o}$, and coarsen $G_{i}$ via the Loop and Clique Coarsening (LCC) to build view $G_{lcc}$. After that, we design the Line Graph Conversion (LGC) to transfer $G_{lcc}$ to view $G_{lgc}$. Subsequently, LCC4GC trains a GT for each view separately to obtain graph-level representation, which is directly concatenated and fed into the downstream classifier.
This chapter is structured as follows: Section 4.2 and Section 4.3 describe LCC and LGC, respectively. Section 4.4 explains the rest of the details of LCC4GC.
4.2 Loop and Clique Coarsening
4.2.1 Algorithm Description
As for coarsening structures into hypernodes, existing methods mainly pay attention to the hierarchical structure of graphs: some cluster nodes based on graph convolution to do graph coarsening [9], and some implement node condensing by iterative sampling [25]. Aggregation of first-order adjacent nodes has intuitive interpretability, which means connections between neighborhoods, while a deeper coarsening shows a poor explanation. Moreover, as mentioned in challenge 3, the particular node disappears into hypernodes, which leads to an ambiguity for topology. Thus, LCC4GC achieves graph coarsening of loops and cliques with shallow coarse-grained clustering restricted by two hard constraints, which magnifies high-level structural information while preserving the characteristic nodes in graphs to the greatest extent. We first give formal definitions of loops and cliques.
Definition 1 (Loop)
Given a graph $G$, a path $G_{l}=\{v_{1}\cdots v_{n+1}\}$ is a loop if there is no repetitive node in $\{v_{1},\cdots,v_{n},v_{n+1}\}$ except $v_{1}=v_{n+1}$ and $(v_{k},v_{k+1})\in\mathcal{E}_{G}$ for $k=1,2,\cdots,n$. $G_{l}$ is a k-loop if $|\mathcal{V}_{G_{l}}|=k$.
Definition 2 (Clique)
A subgraph $G_{c}$ is a clique if $\forall v_{i},v_{j}\in\mathcal{V}_{G_{c}},(v_{i},v_{j})\in\mathcal{E}_{G}(i\neq j)$. $G_{c}$ is a k-clique if $|\mathcal{V}_{G_{c}}|=k$.
Coarsening requires counting all loops and cliques in the graph. However, according to the definitions above, loops are contained within cliques. For example, given a 4-clique marked $Cli=\{ABCD\}$, we can find four 3-loops: $\{ABC\}$, $\{ABD\}$, $\{ACD\}$, and $\{BCD\}$, and one 4-loop $\{ABCD\}$, except for one 4-clique, which is not we want, since identifying $Cli$ as a clique is much more straightforward than five loops. Moreover, since the Maximum Clique Problem (MCP) is NP-hard [42], finding all cliques is also NP-hard. Algorithms require an approximation or a shortcut due to enormous search space.
Mathematically, given an original graph $G_{o}=\{\mathcal{V},\mathcal{E}\}$, LCC aims to rebuild an intermediate coarsening graph $G_{lcc}$ with $n$ nodes where $n\ll|\mathcal{V}|$. In methods [19, 30], supernode $v_{i}$ in $G_{lcc}$ aggregates nodes in $G_{o}$ according to node partitioning sets $\mathcal{P}=\{\mathcal{P}_{1},\cdots,\mathcal{P}_{n}\}$. We use indication matrix $M_{p}\in\{0,1\}^{n\times|\mathcal{V}|}$ to denote the aggregation based on partition set $\mathcal{P}$, where $m_{ij}=1$ if node $j$ in partition $\mathcal{P}_{i}$. We use a new adjacency matrix to represent $G_{lcc}$:
$A_{lcc}=M_{p}AM_{p}^{T},$ (2)
$\bar{A}_{lcc}=diag(a_{ii}^{-1},\cdots,a_{nn}^{-1})(M_{p}AM_{p}^{T}),$ (3)
$H_{lcc}=M_{p}H,$ (4)
$A^{\prime}=A_{L}AA_{L}^{T}=\begin{bmatrix}\mathcal{P}_{1}&E_{12}&\cdots&E_{1n}\\ E_{21}&\mathcal{P}_{2}&\ddots&\vdots\\ \vdots&\ddots&\ddots&\vdots\\ E_{n1}&\cdots&\cdots&\mathcal{P}_{n}\end{bmatrix},$ (5)
$A_{lcc}=\mathcal{I}(A^{\prime}),$ (6)
$\mathcal{I}(X)=\begin{cases}k_{i}&X=\mathcal{P}_{i}\\ 1&\exists x\in X,X=E_{ij},x=1\\ 0&\text{otherwise}\end{cases}.$ (7)
Finally, we give the pseudocode of how the LCC works, as shown in Algorithm 1. Based on Tarjan [43] and Clique Percolation (CP) [44], LCC heuristically iterates the graph hierarchy using loop length and hierarchy depth as constraints for pruning. LCC finds cliques under depths less than $\sigma$ (lines 3-13) first, counting loops only when finding fewer or no cliques (lines 14-21). When updating graphs, LCC reconstructs graphs to build the coarsening view. Figures 2(B) and (C) further depict the workflow of the LCC. We go into details one by one.
Algorithm 1 Loop and Clique Coarsening (LCC)